i need to find the area and perimeter of my room wich is 10x9 feet and take it to inches

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

10*12=120 in
9*12=108 in
P=2*120+2*108
P= 240+216
P=456 in
---------------
A=120*108
A=12,960 sq in

Answer by stanbon(75887)

i need to find the area and perimeter of my room wich is 10x9 feet and take it to inches
--------------
Perimeter = 2(length + width)
P = 2(10+9) ft
P = 38 ft
P = 38(12 inches) = 456 inches
=====================
Cheers,
Stan H.